# Reference.bib

### v1

@@ 8.1. Weakly special subvarieties
Let $(\mathbb{G},D)$ be the Hodge datum associated to the $\mathbb{Z}$-PVHS $\mathbb{V}\rightarrow B$ and the period map $\Phi:B\rightarrow\Gamma\backslash D$. A weak Hodge datum $(\mathbb{H},M)$ is given by a weak Mumford-Tate subdomain $M\leq D$ and its automorphism group $\mathbb{H}\leq\mathbb{G}$ defined over $\mathbb{Q}$. Let $\Gamma_{H}:=\Gamma\cap H:=\mathbb{H_{\mathbb{Q}}}$, we have a natural embedding of Hodge varieties $\Gamma_{H}\backslash M\subset\Gamma\backslash D$.
@@
An irreducible analytic component of $\Phi^{-1}(\Gamma_{H}\backslash M)$ in $B$ is called a weakly special subvariety with Hodge datum $(\mathbb{H},M)$.
A well-known result in Hodge theory and functional transcendence is:
@@
[KO21, Prop. 4.8] $\Phi^{-1}(\Gamma_{H}\backslash M)$ is an algebraic subvariety of $B$. In particular, weakly special subvarieties of $B$ are algebraic.
The proof uses the definability of period mappings [BKT20] and o-minimal geometry [BBT22]. We will give another proof of this result without using o-minimal methods. The main idea is to extend the analytic condition on $B$ defining the weakly special subvarieties to $\overline{B}$ over which a extended period map provided by Theorem 2.8 exists.
@@ 8.2. The period torsors
$\tilde{E}_{\sigma}$ :=$:=\{(q,F)\in\check{E}_{\sigma}\ |\ NF^{\bullet}\subset F^{\bullet-1}\ \forall N\in\sigma(q)\}$
$E_{\sigma}$ :=$:=\{(q,F)\in\tilde{E}_{\sigma}\ |\ (\sigma(q),\mathrm{exp}(\mathrm{log}_{\Gamma_{\sigma}}q)F)\hbox{ is a }\sigma(q)-\hbox{nilpotent orbit}\}$
Here $\sigma(q)\leq\sigma$ is the face corresponds to the torus orbit containing $q$. By [KU08, Chap. 3], $\check{E}_{\sigma}$ admits a natural fs logarithmic analytic structure from $\mathrm{toric}_{\sigma}$ via the projection map, and $\tilde{E}_{\sigma}$ and $E_{\sigma}$ admits the structure of logarithmic manifolds.
By the definition of logarithmic manifolds [KU08, Definition 3.5.7], $\tilde{E}_{\sigma}$ and $E_{\sigma}$ is locally charted by the vanishing loci of some finitely-generated ideal $\mathcal{L}$ of the sheaf of logarithmic $1$-forms $\omega^{1}_{\check{E}_{\sigma}}$ over $\check{E}_{\sigma}$.
By [KU08, Theorem 3.5.10], this ideal sheaf $\mathcal{L}$ locally defining $\tilde{E}_{\sigma}$ and $E_{\sigma}$ is described as follows: Let $\mathcal{F}^{\bullet}$ be the universal Hodge filtration of $\mathcal{O}_{\check{D}}\otimes_{\mathbb{Z}}V_{\mathbb{C}}$ over $\check{D}$, $\mathcal{G}:=\mathcal{O}_{\check{D}}\otimes_{\mathbb{C}}\sigma_{\mathbb{C}}$ and $\mathcal{G}^{-1}:=\{X\in\mathcal{G},\ X\mathcal{F}^{\bullet}\subset\mathcal{F}^{\bullet-1}\}$. Let $\mathcal{P}$ be the pullback of $\mathcal{G}/\mathcal{G}^{-1}$ to $\tilde{E}_{\sigma}$ via the projection map. There is a canonical projection morphism $\eta:\mathcal{O}_{\check{E}_{\sigma}}\otimes_{\mathbb{C}}\sigma_{\mathbb{C}}\rightarrow\mathcal{P}$ giving a global section of $(\mathcal{O}_{\check{E}_{\sigma}}\otimes_{\mathbb{C}}\sigma_{\mathbb{C}})^{\vee}\otimes\mathcal{P}$ which is isomorphic to $\omega^{1}_{\check{E}_{\sigma}}\otimes\mathcal{P}$. Therefore, the ideal $\mathcal{L}\subset\omega^{1}_{\check{E}_{\sigma}}$ is generated by the coordinates of $\eta$ under local frames of $\mathcal{P}$.
@@
(8.5) $a\in\sigma_{\mathbb{C}}:E_{\sigma}\rightarrow E_{\sigma},\ (a,(q,F))\rightarrow(\mathbb{e}(a)q,\exp(-a)F)$
where the map $\mathbb{e}(a)$ is defined by
(8.6) $\mathbb{e}(z\log(\gamma))=\exp(2\pi iz)\otimes\gamma\in\mathrm{torus}_{\sigma}.$
This realizes $E_{\sigma}$ as a $\sigma_{\mathbb{C}}$-torsor over $\Gamma_{\sigma}\backslash D_{\sigma}$ in the category of logarithmic manifolds by [KU08, Section 7.2].
According to [KU08, Section 7.3.5], for $x:=(q,F)\in E_{\sigma}\subset\check{E}_{\sigma}$, the structure of logarithmic manifold on $\Gamma_{\sigma}\backslash D_{\sigma}$ around $\Theta_{\sigma}(x)$ is described as follows: Let $\check{A}_{x}\subset\check{E}_{\sigma}$ be an analytic variety through $x$ such that:
@@
Let $A_{x}:=\check{A}_{x}\cap E_{\sigma}$, then $A_{x}\subset E_{\sigma}$ acquires a structure of logarithmic submanifold of $E_{\sigma}$, and by [KU08, Lemma 7.3.3], We may associate the logarithmic manifold structure on $\Gamma_{\sigma}\backslash D_{\sigma}$ by regarding $A_{x}$ as a chart around $\Theta_{\sigma}(x)$.
@@ 8.3. The relative period torsors
For the fixed weak Hodge datum $(\mathbb{H},M)$ of $(\mathbb{G},D)$. Denote $\mathfrak{h}:=\mathrm{Lie}(\mathbb{H})$ which is a Lie subalgebra of $\mathfrak{g}$, and $\sigma_{\mathfrak{h}}:=\sigma\cap\mathfrak{h}$. Define
@@
where $\check{M}$ is the compact dual of $M$. $\check{E}_{\sigma,M}$ may be realized as an algebraic subvariety of $\check{E}_{\sigma}$ via the embedding of cones $\sigma_{\mathfrak{h}}\hookrightarrow\sigma$. Let $\tilde{E}_{\sigma,M}:=\tilde{E}_{\sigma}\cap\check{E}_{\sigma,M}$ and $E_{\sigma,M}:=E_{\sigma}\cap\check{E}_{\sigma,M}$. Endow each space with the subspace topology.
Endow $\check{E}_{\sigma,M}$ the fs logarithmic analytic structure from restricting the one on $\check{E}_{\sigma,M}$. This is different from the structure on $\check{E}_{\sigma,M}$ as an independent period torsor. Let $\mathfrak{M}:=\Theta_{\sigma}(E_{\sigma,M})\subset\Gamma_{\sigma}\backslash D_{\sigma}$. As a consequence, $\tilde{E}_{\sigma,M}$ (resp. $E_{\sigma,M}$) is a logarithmic submanifold of $\tilde{E}_{\sigma}$ (resp. $E_{\sigma}$). Denote $\Theta_{\sigma,M}$ as the restriction of $\Theta_{\sigma}$ on $E_{\sigma,M}$.
@@
The map $\Theta_{\sigma,M}:E_{\sigma,M}\rightarrow\mathfrak{M}$ is a $\sigma_{\mathfrak{h},\mathbb{C}}$-torsor of logarithmic manifolds. In particular, $\mathfrak{M}$ is a logarithmic submanifold of $\Gamma_{\sigma}\backslash D_{\sigma}$.
@@
We first show $\sigma_{\mathfrak{h},\mathbb{C}}$ acts freely on the fiber of $\Theta_{\sigma,M}$. From the result for $\Theta_{\sigma}:E_{\sigma}\rightarrow\Gamma_{\sigma}\backslash D_{\sigma}$, it is enough to show on the set-theoretical level that for $q\in\mathfrak{M}$, the fiber $\Theta^{-1}_{\sigma,M}(q)\subset E_{\sigma,M}$ is a $\sigma_{\mathfrak{h},\mathbb{C}}$-orbit. Fix $(q,F)\in E_{\sigma,M}$, for any $a\in\sigma_{\mathbb{C}}$, by Equation (8.5) it is enough to show $(\mathbb{e}(a)q,\exp(-a)F)\in E_{\sigma,M}$ if and only if $a\in\sigma_{\mathfrak{h},\mathbb{C}}$. For the first factor, note that $\mathbb{e}(a)q\in\mathrm{toric}_{\sigma_{\mathfrak{h}}}$ if and only if $\mathbb{e}(a)\in\mathrm{torus}_{\sigma_{\mathfrak{h}}}\simeq C^{*}\otimes\Gamma_{\sigma_{\mathfrak{h}}}$ if and only if $a\in\sigma_{\mathfrak{h}}$. For the second factor, note that the natural map $\sigma_{\mathbb{C}}\rightarrow T_{\check{D}}F$ is an injection, hence $\exp(-a)F\in M$ if and only if $a\in T_{\check{M}}F\cap\sigma_{\mathbb{C}}=\sigma_{\mathfrak{h},\mathbb{C}}$.
Next we show the properness of the $\sigma_{\mathfrak{h},\mathbb{C}}$-action. By [KU08, Lemma 7.2.7], it is enough to show for any directed families $\{(q_{\lambda},F_{\lambda})\}_{\lambda\in\Lambda}\subset E_{\sigma,M}$ and $\{a_{\lambda}\}_{\lambda\in\Lambda}\subset\sigma_{\mathfrak{h},\mathbb{C}}$ such that both $\{(q_{\lambda},F_{\lambda})\}_{\lambda\in\Lambda}$ and $\{a_{\lambda}(q_{\lambda},F_{\lambda})\}_{\lambda\in\Lambda}$ converge in $E_{\sigma,M}$, $\{a_{\lambda}\}$ must converge in $\sigma_{\mathfrak{h},\mathbb{C}}$. This immediately follows from the facts that $\{a_{\lambda}\}$ converges in $\sigma_{\mathbb{C}}$ because of the proper action of $\sigma_{\mathbb{C}}$ on $E_{\sigma}$, and $\sigma_{\mathfrak{h},\mathbb{C}}$ is a closed subgroup of $\sigma_{\mathbb{C}}$.
Finally, similar to the ideas used in [KU08, Section 7.3.5], we may choose $x=(q,F)\in\check{B}_{x}\subset\check{E}_{\sigma,M}$ such that $\check{B}_{x}\simeq B_{1}\times B_{2}$, where $B_{1}$ is an open neighborhood of $q$ in $\mathrm{toric}_{\sigma(q)\cap\mathfrak{h}}$, and $F\in B_{2}\subset M$ intersects $\exp((\sigma(q)\cap\mathfrak{h})_{\mathbb{C}})F$ transversely at $F$ and with complementary dimension in $\check{M}$. Let $B_{x}:=\tilde{B}_{x}\cap E_{\sigma,M}$. From the arguments we have in the last section for $\tilde{A}_{x}$ and $A_{x}$, this $B_{x}\subset E_{\sigma,M}$ satisfies the requirement. More precisely, the inclusions $\check{B}_{x}\subset\check{A}_{x}$ and $B_{x}\subset A_{x}$ realize $\mathfrak{M}$ as a logarithmic submanifold of $\Gamma_{\sigma}\backslash D_{\sigma}$. The proof is complete. ∎
@@ 8.4. Proof of the algebraicity result
We are ready to prove Theorem 8.2. Fix a completion $\overline{\Phi}:\overline{B}\rightarrow\overline{\wp}$ given by Theorem 2.8. Once we can show any weakly special subvariety $S\subset B$ with Hodge datum $(\mathbb{H},M)$ can be extended to an analytic subvariety of $\overline{B}$, Serre’s GAGA theorem will grant the algebraicity of $S$.
Take the open cover $\overline{B}=\bigcup_{k}W_{k}$ given by Theorem 5.1. The restriction of $\overline{\Phi}$ on each $W_{k}$ can be lifted to the local KNU maps $\overline{\Phi}_{k}:W_{k}\rightarrow\Gamma_{\tau_{k}}\backslash D_{\tau_{k}}$ for some monodromy cone $\tau_{k}\subset\mathfrak{g}$. Theorem 3.12 and 5.1 implies each $\overline{\Phi}_{k}(W_{k})$ is a complex analytic space.
Let $\mathfrak{M}_{k}$ be the image of $(\mathrm{toric}_{\tau_{k}\cap\mathfrak{h}}\times M)\cap E_{\tau_{k},M}$ in $\Gamma_{\tau_{k}}\backslash D_{\tau_{k}}$ via the torsor map (8.4). Let $x\in\mathfrak{M}_{k}\cap\overline{\Phi}_{k}(W_{k})$. By Proposition 8.4, we may choose neighborhoods $x\in A_{x}\subset\Gamma_{\tau_{k}}\backslash D_{\tau_{k}}$, $x\in V_{x}\subset\overline{\Phi}_{k}(W_{k})$ and $B_{x}:=A_{x}\cap\mathfrak{M}_{k}$ such that $V_{x}\cap B_{x}$ is an analytic subspace of $V_{x}\cap A_{x}$. This implies for each $k$, $\overline{\Phi}_{k}^{-1}(\mathfrak{M}_{k})$ is an analytic subspace of $W_{k}$.
Finally, since each $\overline{\Phi}_{k}$ has connected fibers, we may patch all $\overline{\Phi}_{k}^{-1}(\mathfrak{M}_{k})$ to an analytic subvariety $S$ of $\overline{B}$ and hence must be algebraic by GAGA, and it is evident that $S\cap B=\Phi^{-1}(\Gamma_{H}\backslash M)$. The proof of Theorem 8.2 is complete.

### v4

@@ 7.1. Weakly special subvarieties
Let $(\mathbf{G},D)$ be the Hodge datum associated to the $\mathbb{Z}$-PVHS $\mathbb{V}\rightarrow B$ and the period map $\D\Phi:B\rightarrow\Gamma\backslash$ D. A weak Hodge datum $(\mathbf{H},M)$ is given by a weak Mumford-Tate subdomain $M\leq D$ and its automorphism group $\mathbf{H}\leq\mathbf{G}$ defined over $\mathbb{Q}$. Let $\Gamma_{H}:=\Gamma\cap H:=\mathbf{H_{\mathbb{Q}}}$, we have a natural embedding of Hodge varieties $\Gamma_{H}\backslash M\subset\Gamma\backslash D$.
@@
[…]
As mentioned above, the following is a well-known result in Hodge theory and functional transcendence. It is first proved by [CDK95] for the case of special subvarieties, then generalized to the weakly case.
@@
[…]
The proof uses the definability of period mappings [BKT20] and o-minimal geometry [BBT22]. We will give another proof of this result using the generalized toroidal completion given by Theorem 2.8. The main idea is to extend the analytic condition on $B$ defining the weakly special subvarieties to $\overline{B}$ over which a extended period map provided by Theorem 2.8 exists. This may be seen as a generalization of [CDK95] using the language of logarithmic geometry.
@@ 7.2. The period torsors
$\tilde{E}_{\sigma}$ $:=\{(q,F)\in\check{E}_{\sigma}\ |\ NF^{\bullet}\subset F^{\bullet-1}\ \forall N\in\sigma(q)\}$
[…]
Here $\sigma(q)\leq\sigma$ is the face corresponds to the torus orbit containing $q$.
Following Kato–Usui’s book [KU08], we will see these spaces carry rich structures in the theory of logarithmic geometry. Regarding basic definitions and examples in logarithmic geometry including logarithmic analytic spaces, logarithmic differential forms, and the strong topology, we refer readers to [KU08, Chap. 2,3]. Here we introduce one of the most important definitions.
Definition 7.3.
[KU08, Def. 3.5.7] By a logarithmic manifold $X$, we mean a $\mathbb{C}$-logarithmic local ringed space with an open cover $\{U_{\lambda}\}_{\lambda\in\Lambda}$ such that for each $\lambda\in\Lambda$ there exists a logarithmically smooth fs logarithmic analytic space $Z_{\lambda}$ and a finite subset $I_{\lambda}$ of logarithmic differential $1$-forms on $Z_{\lambda}$, such that
$U_{\lambda}\simeq\{z\in Z_{\lambda},\ I_{\lambda,z}=0\}$
where the RHS is endowed with the strong topology in $Z_{\lambda}$ and the pull back of the logarithmic structure on $Z_{\lambda}$. In this case, we call $(Z_{\lambda},I_{\lambda})$ a chart of $X$.
By [KU08, Chap. 3], $\check{E}_{\sigma}$ admits a natural fs logarithmic analytic structure from $\mathrm{toric}_{\sigma}$ via the projection map, and $\tilde{E}_{\sigma}$ and $E_{\sigma}$ admits the structure of logarithmic manifolds. Their local charts are described as follows: Let $\mathcal{F}^{\bullet}$ be the universal Hodge filtration of $\mathcal{O}_{\check{D}}\otimes_{\mathbb{Z}}V_{\mathbb{C}}$ over $\check{D}$, $\mathcal{G}:=\mathcal{O}_{\check{D}}\otimes_{\mathbb{C}}\sigma_{\mathbb{C}}$ and $\mathcal{G}^{-1}:=\{X\in\mathcal{G},\ X\mathcal{F}^{\bullet}\subset\mathcal{F}^{\bullet-1}\}$. Let $\mathcal{P}$ be the pullback of $\mathcal{G}/\mathcal{G}^{-1}$ to $\tilde{E}_{\sigma}$ via the projection map. There is a canonical projection morphism $\eta:\mathcal{O}_{\check{E}_{\sigma}}\otimes_{\mathbb{C}}\sigma_{\mathbb{C}}\rightarrow\mathcal{P}$ giving a global section of $(\mathcal{O}_{\check{E}_{\sigma}}\otimes_{\mathbb{C}}\sigma_{\mathbb{C}})^{\vee}\otimes\mathcal{P}$ which is isomorphic to $\omega^{1}_{\check{E}_{\sigma}}\otimes\mathcal{P}$. Therefore, the ideal $\mathcal{L}\subset\omega^{1}_{\check{E}_{\sigma}}$ is generated by the coordinates of $\eta$ under local frames of $\mathcal{P}$.
@@
(7.5) $a\in\sigma_{\mathbb{C}}:E_{\sigma}\rightarrow E_{\sigma},\ (a,(q,F))\rightarrow(\mathbf{e}(a)q,\exp(-a)F)$
where the map $\mathbf{e}(a)$ is defined by
(7.6) $\mathbf{e}(z\log(\gamma))=\exp(2\pi iz)\otimes\gamma\in\mathrm{torus}_{\sigma}.$
[…]
@@
[…]
@@ 7.3. The relative period torsors
For the fixed weak Hodge datum $(\mathbf{H},M)$ of $(\mathbf{G},D)$. Denote $\mathfrak{h}:=\mathrm{Lie}(\mathbf{H})$ which is a Lie subalgebra of $\mathfrak{g}$, and $\sigma_{\mathfrak{h}}:=\sigma\cap\mathfrak{h}$. Define
@@
where $\check{M}$ is the compact dual of $M$. $\check{E}_{\sigma,M}$ may be realized as an algebraic subvariety of $\check{E}_{\sigma}$ via the embedding of cones $\sigma_{\mathfrak{h}}\hookrightarrow\sigma$. Let $\tilde{E}_{\sigma,M}:=\tilde{E}_{\sigma}\cap\check{E}_{\sigma,M}$ and $E_{\sigma,M}:=E_{\sigma}\cap\check{E}_{\sigma,M}$. Endow each space with the subspace topology. Endow $\check{E}_{\sigma,M}$ the fs logarithmic analytic structure from restricting the one on $\check{E}_{\sigma,M}$. This is different from the structure on $\check{E}_{\sigma,M}$ as an independent period torsor.
Definition 7.5.
Let $X$ be a logarithmic manifold. We say $Y\subset X$ is a logarithmic submanifold if it is a logarithmic manifold as well as a $\mathbb{C}$-logarithmic local ringed subspace of $X$, and for any $y\in Y\subset X$, there exists a local chart $(Z_{\lambda},I_{\lambda})$ of $X$ around $x$ such that there is a logarithmically smooth fs logarithmic analytic subspace $Z_{\lambda}^{\prime}\leq Z_{\lambda}$, and $(Z_{\lambda}^{\prime},I_{\lambda}|_{Z_{\lambda}^{\prime}})$ is a local chart of $y$ in $Y$.
Let $\Theta_{\sigma,M}$ be the restriction of $\Theta_{\sigma}$ on $E_{\sigma,M}$ and $\mathfrak{M}$ be its image. It is clear that $\tilde{E}_{\sigma,M}$ (resp. $E_{\sigma,M}$) is a logarithmic submanifold of $\tilde{E}_{\sigma}$ (resp. $E_{\sigma}$). We will show:
@@
[…]
@@
[…]
Next we show the properness of the $\sigma_{\mathfrak{h},\mathbb{C}}$-action. By [KU08, Lemma 7.2.7], it is enough to show for any directed families $\{(q_{\lambda},F_{\lambda})\}_{\lambda\in\Lambda}\subset E_{\sigma,M}$ and $\{a_{\lambda}\}_{\lambda\in\Lambda}\subset\sigma_{\mathfrak{h},\mathbb{C}}$ such that both $\{(q_{\lambda},F_{\lambda})\}_{\lambda\in\Lambda}$ and $\{a_{\lambda}(q_{\lambda},F_{\lambda})\}_{\lambda\in\Lambda}$ converge in $E_{\sigma,M}$, $\{a_{\lambda}\}$ must converge in $\sigma_{\mathfrak{h},\mathbb{C}}$. This immediately follows from the facts that $\{a_{\lambda}\}$ converges in $\sigma_{\mathbb{C}}$ because of the proper action of $\sigma_{\mathbb{C}}$ on $E_{\sigma}$, and $\sigma_{\mathfrak{h},\mathbb{C}}$ is a closed subspace of $\sigma_{\mathbb{C}}$.
[…]
@@ 7.4. Proof of the algebraicity result
We are ready to prove Theorem 7.2. Fix a completion $\overline{\Phi}:\overline{B}\rightarrow\overline{\wp}$ given by Theorem 2.8. Once we can show any weakly special subvariety $S\subset B$ with Hodge datum $(\mathbf{H},M)$ can be extended to an analytic subvariety of $\overline{B}$, Serre’s GAGA theorem will grant the algebraicity of $S$.